Scientist in Alaska recently found a hadrosaur footprint that measured 14 inches across. It is believed that the footprint was created by a young dinosaur that was approximately 27 feet long. Adult hadrosaurs have been known to be 40 feet long. How wide would you expect a footprint of an adult hadrosaur to be? show your reasoning.
Answer by stanbon(75887)

Scientist in Alaska recently found a hadrosaur footprint that measured 14 inches across. It is believed that the footprint was created by a young dinosaur that was approximately 27 feet long. Adult hadrosaurs have been known to be 40 feet long. How wide would you expect a footprint of an adult hadrosaur to be? show your reasoning.
---------
Use a proportion.
---
x/40 = 14/27
x = 40*14/27
x = 20.47 inches
=========================
Cheers,
Stan H.
=========================